OpenStudy (anonymous):

1/(1+i) - 1/(1-i) please help!

OpenStudy (welshfella):

first what is the LCM of 1-i and 1+i?

OpenStudy (anonymous):

2

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

2 = 1 - i^2 now multiply both fractions by this to get the numerator

OpenStudy (welshfella):

= 1 - 1 - (1 + i) ------------ 2

OpenStudy (anonymous):

is the answer -i?

OpenStudy (welshfella):

1 - i - (1 + i) ------------ 2 = -2i / 2 = -i

OpenStudy (anonymous):

thank you so much!

OpenStudy (welshfella):

yw